# for Large Language Models

### v3

@@ A.6 API Services Available in EasyInstruct
GPT-3.5
A set of models that improve on GPT-3 and can understand as well as generate natural language or code.
gpt-3.5-turbo
GPT-4
A set of models that improve on GPT-3.5 and can understand as well as generate natural language or code.
gpt-4
@@
Claude
A next-generation AI assistant based on Anthropic’s research into training helpful, honest, and harmless AI systems.
claude-2
Claude-Instant
A lighter, less expensive, and much faster option than Claude.
claude-instant-1
@@
Command
An instruction-following conversational model that performs language tasks with high quality, more reliably, and with a longer context than cohere’s base generative models.
command
Command-Light
A smaller, faster version of Command. Almost as capable, but a lot faster.
command-light
@@
A.7 Example features in the Prompts module

### v4

@@ A.5 API Services Available in EasyInstruct
GPT-3.5 A set of models that improve on GPT-3 and can understand as well as generate natural language or code. gpt-3.5-turbo
GPT-4 A set of models that improve on GPT-3.5 and can understand as well as generate natural language or code. gpt-4
@@
Claude A next-generation AI assistant based on Anthropic’s research into training helpful, honest, and harmless AI systems. claude-2
Claude-Instant A lighter, less expensive, and much faster option than Claude. claude-instant-1
@@
Command An instruction-following conversational model that performs language tasks with high quality, more reliably, and with a longer context than cohere’s base generative models. command
Command-Light A smaller, faster version of Command. Almost as capable, but a lot faster. command-light
@@
A.6 Example features in the Prompts module